# LumiMAS: A Comprehensive Framework for Real-Time Monitoring and Enhanced Observability in Multi-Agent Systems

### v1

@@ 4.2 Anomaly Detection Layer
Combined Latent Space AE.
@@
Log Classification.
During inference, each execution log’s features are extracted and propagated through the model, and its reconstruction error is computed to determine whether the log is anomalous or benign. The classification threshold can be selected by optimizing a performance metric (e.g., the F1 score). The logs predicted as anomalous by the anomaly detection layer raise an alert to the LumiMAS user and are forwarded to the anomaly explanation layer.
4.3 Anomaly Explanation Layer
While the alerts produced by the anomaly detection layer provide an initial indication of abnormal system behaviors, LumiMAS extends this capability by integrating an LMA classification agent guided by a tailored prompt.
@@ 4.3 Anomaly Explanation Layer
The classification agent analyzes the suspected log, aiming to assign the appropriate category out of a predefined set of vulnerability types (e.g., bias, DPI). Moreover, by incorporating the ’benign’ category, the agent can filter cases that do not indicate anomalous behavior, which can reduce false positives. The classification agent’s prompt, which is provided in Appendix B, is composed of four structured elements: (1) agent’s role, (2) agent’s task description that outlines expected behavior, (3) a list of the defined vulnerabilities, and (4) format of expected response.
Further deepening the insights, a dedicated RCA LMA is employed to locate the failure source, utilizing the vulnerability type predicted by the classification agent. The RCA agent is driven by a prompt designed similarly to the classification agent, allowing it to analyze the chronological execution flow, including inter-agent interactions. The agent’s output includes an explanation for its prediction, providing a detailed account of the failure’s root cause.
Working collaboratively, the classification and RCA agents form a cohesive diagnostic team. Together, they provide a holistic analysis: first, they characterize the nature of each anomaly and then locate its origin within the system’s execution flow.
5 Evaluation
5.1 Evaluation Settings
@@
This subsection outlines the primary settings of our evaluation. Additional details, such as extended hyperparameter configurations and information about the applications used, are provided in the Appendix C.
MAS Applications.
@@
We employed seven different MAS applications, each demonstrating a distinct vulnerability, and two MAS platforms, CrewAI (CrewAI Inc. 2025) and LangGraph (LangGraph 2023). On the CrewAI framework, we employed the Trip Planner and Instagram Post applications (Crew AI Inc. 2024) to demonstrate the detection of DPI and IPI attacks, respectively. For MP attack detection, we utilized the Real Estate Team application (Hancock 2024), which includes an RAG database. On the LangGraph framework, we employed the GenFic application and our adapted version of the Trip Planner application from (Crew AI Inc. 2024), focusing on DPI attack detection.
@@
While some vulnerabilities are more straightforward to generate and evaluate (e.g., DPI and IPI), vulnerabilities that originate from inherent limitations of the LLM (e.g., hallucination and bias) are more difficult to trigger. Therefore, we developed FTAs to enable the assessment of detection and mitigation strategies. HalluCheck, our hallucination-tailored application, utilizes a question-answering dataset (Rajpurkar et al. 2016) with three specially designed agents to illustrate the effect on the features (e.g., increased run time) in a MAS application when a hallucination occurs. Similarly, BiasCheck enables illustrating the effect on the features in a MAS when bias occurs, leveraging a stereotype questions dataset (Parrish et al. 2021).
Failure Generation.
While the failures in our FTAs result from the application’s design and datasets used, the faults in the other applications are induced in different ways, aligned with the type of vulnerability.
@@
The generation of DPI, IPI, and MP attacks was inspired by (Deng et al. 2025; Zhang et al. 2024b) and allowed us to generate realistic attack scenarios. In the first DPI scenario, we crafted a prompt that forces the agent’s LLM to produce outputs in an invalid format. As a result, the agent fails to parse the response, which leads to increased computational load, excessive token usage, and longer execution time. In the second DPI, we concatenated misinformation with the user input, causing the agent to base its reasoning and outputs on false information aligned with the attacker’s intent. Another DPI variant acted as a form of backdoor attack, injecting into the prompt condition tied to the agent’s identity (e.g., name). Once triggered, the agent’s behavior changes to serve the attacker’s intent and affects the system’s overall output. To demonstrate the IPI scenario, we set up a malicious web server containing HTML content with false instructions. This attack aimed to lure the agent into repeatedly following deceptive URLs, resulting in another form of system exhaustion. Lastly, in the MP scenario, we poisoned the RAG database that the system’s agents rely on, which results in poisoned information being retrieved.
Data Collection and Splitting.
@@
To demonstrate the robustness of our framework, we evaluate its effectiveness using two different LLMs as the underlying models for the agents: OpenAI GPT-4o mini (Hurst et al. 2024) and OpenAI o3-mini [^1] https://openai.com/index/openai-o3-mini/. GPT-4o mini is a lightweight version of GPT-4o, optimized for speed and efficiency, with strong general performance. o3-mini is a compact LLM that has been shown to perform efficiently across various natural language processing tasks.
Vulnerability
Method Performance Overhead
Accuracy $\uparrow$ F1 $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ FPR $\downarrow$ Latency $\downarrow$ Cost $\downarrow$
Hallucination Consistency of 3 0.700 0.727 0.800 0.667 0.400 9.215 5189.425
Consistency of 5 0.695 0.716 0.770 0.670 0.380 15.676 8648.480
LLM-as-a-judge 0.595 0.382 0.250 0.806 0.060 7.138 2516.050
Agent-as-a-judge 0.595 0.409 0.280 0.757 0.090 12.312 1341349.905
EPI (Ours) 0.745 0.730 0.690 0.775 0.200 0.003 -
Semantic (Ours) 0.720 0.682 0.600 0.790 0.160 0.014 -
Combined (Ours) 0.765 0.728 0.630 0.863 0.100 0.017 -
Bias toxic-bert 0.510 0.671 1.000 0.505 0.980 0.398 -
LLM-as-a-judge 0.905 0.907 0.930 0.886 0.120 7.810 2324.075
Agent-as-a-judge 0.900 0.900 0.900 0.900 0.100 14.042 1283347.355
EPI (Ours) 0.625 0.675 0.780 0.595 0.530 0.016 -
Semantic (Ours) 0.600 0.615 0.640 0.593 0.440 0.020 -
Combined (Ours) 0.660 0.670 0.690 0.651 0.370 0.024 -
Direct
Prompt
Injection (DPI)
LLM-as-a-judge 0.605 0.696 0.913 0.563 0.703 9.759 4503.273
Agent-as-a-judge 0.705 0.742 0.870 0.648 0.460 25.867 4783762.972
EPI (Ours) 0.742 0.765 0.873 0.685 0.390 0.008 -
Semantic (Ours) 0.770 0.794 0.900 0.712 0.360 0.074 -
Combined (Ours) 0.767 0.761 0.803 0.737 0.270 0.088 -
Indirect
Prompt
Injection (IPI)
LLM-as-a-judge 0.560 0.679 0.930 0.534 0.810 20.182 6844.650
Agent-as-a-judge 0.650 0.724 0.920 0.597 0.620 29.617 9929870.045
EPI (Ours) 0.965 0.966 1.000 0.935 0.070 0.029 -
Semantic (Ours) 0.890 0.894 0.930 0.861 0.150 0.102 -
Combined (Ours) 0.970 0.970 0.970 0.970 0.030 0.142 -
Memory
Poisoning (MP)
LLM-as-a-judge 0.805 0.820 0.890 0.761 0.280 9.509 3764.930
Agent-as-a-judge 0.780 0.784 0.800 0.769 0.240 13.348 1787606.550
EPI (Ours) 0.540 0.643 0.830 0.525 0.750 0.003 -
Semantic (Ours) 0.690 0.710 0.760 0.667 0.380 0.028 -
Combined (Ours) 0.500 0.561 0.640 0.500 0.640 0.032 -
@@
Table 1: Anomaly detection results obtained using the CrewAI apps, with GPT-4o-mini as the underlying model.
Baselines.
@@
As no prior work had been done to detect failures in MASs, we utilized the LLM-as-a-judge paradigm (Zheng et al. 2023) with an engineered prompt for the anomaly detection task. Inspired by Zhuge et al. (2024), we also implemented the agent-as-a-judge variant, with the task and agent definition based on the same engineered prompt. We equip the agent with scraping and searching tools, enabling it to retrieve additional information to improve its effectiveness in detecting failures. GPT-4o (Hurst et al. 2024) was used as the backbone of our baselines, analyzing each agent’s input and output. Consequently, the prediction of the baseline approaches also indicate which agent is responsible for the failure, and this information is used in our RCA evaluation. Additional results using the smaller and more cost-efficient GPT-4o-mini (Hurst et al. 2024) are provided in Appendix C.
We also evaluate existing hallucination and bias detection strategies for our FTA apps. For hallucination detection, we adopted the consistency comparison approach (Manakul, Liusie, and Gales 2023), in which each agent was instructed to repeat the task 3 or 5 additional times, and the original result was considered consistent (i.e., benign) if it aligned with the majority of the outputs. Regarding bias, we employed toxic-bert (Hanu and Unitary team 2020) to assess the bias in each corresponding agent.
Metrics.
@@
To evaluate failure detection performance, we used several standard metrics applied in binary classification tasks (accuracy, precision, recall, F1 score, and false positive rate (FPR)), as well as metrics measuring the solutions’ overhead resource consumption (average inference time and token count (Cost) for models that require payment). For the RCA evaluation, we used the root cause accuracy (RA) metric from (Zhang et al. 2024c; Zhou et al. 2023). In the results tables, the best performance in each column is marked in bold, and the second-best is underlined. The arrows ($\uparrow$, $\downarrow$) indicate whether the optimal value is higher or lower.
Implementation Details.
@@
We implemented our method using Python 3.11 and PyTorch 2.6, and the models were trained on an RTX 4090 GPU. Each application had its own set of hyperparameters, and all models were trained using the Adam optimizer. For the hallucination, the EPI AE model was configured with a single LSTM layer and trained for 150 epochs using a learning rate of 1.2e-3 and a batch size of 16. For the semantic AE, the textual features (i.e., the agent’s LLM interactions) were embedded using the all-MiniLM-L6-v2 model based on sentence-transformers (Reimers and Gurevych 2019). A bidirectional LSTM architecture was used in the semantic AE model, which was trained with a learning rate of 8.3e-4 for 150 epochs and a batch size of 64. The combined AE was trained for 260 epochs with a learning rate of 1.5e-4 and a batch size of 16. The anomaly detection classification threshold was selected based on the F1 score on the validation set, while the models’ other hyperparameters were tuned to minimize the reconstruction error on the benign validation set. The distance function used in all three models’ loss function was the mean square error.
5.2 Evaluation Results

### v2

@@ 4.2.2. Model Training
4.2.3. Combined Latent Space AE
@@
4.2.4. Failure-Detector Mapping
Each LumiMAS’s detector aligns with different signal types and corresponding failure categories within the MAST taxonomy cemri2025multi. The Semantic-AE operates over embeddings of agent reasoning, message exchanges, and memory traces, enabling it to capture linguistic and logical inconsistencies, reasoning drift, and other deviations from prompt specifications or agent communication. These correspond to MAST’s Specification Issues and Inter-Agent Misalignment failures, such as Disobey Task Specification, Disobey Role Specification, Loss of Conversation History, Task Derailment, Conversation Reset, and Reasoning–Action Mismatch. The EPI-AE, in contrast, consumes execution and performance indicators, including latency, message frequency, and runtime metrics, and is thus sensitive to abnormal execution patterns, resource exhaustion, and unexpected termination. These signals are characteristic of MAST’s failures, including Premature Termination, Conversation Reset, and Step Repetition. Finally, the combined detector integrates both semantic and EPI features to uncover inconsistencies between an agent’s communicative intent and its observable behavior. This fusion provides greater robustness across the anomaly types. Nevertheless, certain failure modes still benefit more from specialized components when the anomaly is strongly semantic or operation-oriented. Together, these detectors provide complementary coverage of the major MAST failure categories. While LumiMAS achieves broad coverage across failure types, its current architecture is not primarily focused on Task Verification failures (categories 2.2, 3.2, and 3.3 in MAST), which typically require explicit reasoning about task completeness or output validation. As a result, failures such as Incomplete Verification or Incorrect Verification may be indirectly detected through performance anomalies (e.g., token count or latency changes) but are not explicitly modeled.
4.3. Anomaly Explanation Layer
While producing real-time alerts for incoming threats is the primary goal of our framework, providing additional contextual information alongside each detection can help to effectively address the anomaly. Aiming to provide additional context to the alerts rising from the anomaly detection layer LumiMAS features an anomaly explanation layer which incorporate two specialized LMA agents: (1) classification agent and (2) root cause analysis (RCA) agent.
@@ 4.3. Anomaly Explanation Layer
The classification agent analyzes the suspected log, aiming to assign the appropriate category out of a predefined set of vulnerability types (e.g., bias, DPI). Moreover, by incorporating the ’benign’ category, the agent can filter cases that do not indicate anomalous behavior, which can reduce false positives. The classification agent’s prompt, which is provided in Appendix B, is composed of four structured elements: (1) the agent’s role, (2) the agent’s task description that outlines expected behavior, (3) a list of the defined vulnerabilities, and (4) the format of expected response.
The RCA agent leverages the classification output to trace the origin of the detected anomaly. Guided by a similarly structured prompt, the RCA agent examines the chronological execution flow and inter-agent interactions to identify the first agent responsible for introducing the anomalous behavior. Its output includes both the root cause agent identifier and a detailed explanation of the causal chain leading to the failure. Working collaboratively, the classification and RCA agents form a cohesive diagnostic pipeline: the first characterizes the nature of the anomaly, and the second locates its origin within the system’s execution flow, enabling comprehensive and interpretable MAS failure analysis.
5. Evaluation
5.1. Evaluation Settings
@@
This subsection outlines the primary settings of our evaluation. Additional details, such as extended hyperparameter configurations and information about the applications used, are provided in Appendix C.
5.1.1. MAS Applications
@@
We employed seven different MAS applications, each demonstrating a distinct vulnerability, and two MAS platforms, CrewAI crewai and LangGraph LangGraphMisc. On the CrewAI framework, we employed the Trip Planner and Instagram Post applications crewAI-examples to demonstrate the detection of DPI and IPI attacks, respectively. For MP attack detection, we utilized the Real Estate Team application crewai-rag-deep-dive, which includes an RAG database. On the LangGraph framework, we employed the GenFic application and our adapted version of the Trip Planner application from crewAI-examples, focusing on DPI attack detection.
@@
While some vulnerabilities are more straightforward to generate and evaluate (e.g., DPI and IPI), vulnerabilities that originate from inherent limitations of the LLM (e.g., hallucination and bias) are more difficult to trigger. Therefore, we developed FTAs to enable the assessment of detection and mitigation strategies. HalluCheck, our hallucination-tailored application, utilizes a question-answering dataset rajpurkar2016squad with three specially designed agents to illustrate the effect on the features (e.g., increased run time) in a MAS application when a hallucination occurs. Similarly, BiasCheck enables illustrating the effect on the features in a MAS when bias occurs, leveraging a stereotype questions dataset parrish2022bbq.
5.1.2. Failure Generation
@@
While the failures in our FTAs result from the application’s design and datasets used, the faults in the other applications are induced in different ways, aligned with the type of vulnerability. The generation of DPI, IPI, and MP attacks was inspired by deng2025ai; zhang2024agent and allowed us to generate realistic attack scenarios. In the first DPI scenario, we crafted a prompt that forces the agent’s LLM to produce outputs in an invalid format. As a result, the agent fails to parse the response, which leads to increased computational load, excessive token usage, and longer execution time. In the second DPI, we concatenated misinformation with the user input, causing the agent to base its reasoning and outputs on false information aligned with the attacker’s intent. Another DPI variant acted as a form of backdoor attack, injecting into the prompt condition tied to the agent’s identity (e.g., name). Once triggered, the agent’s behavior changes to serve the attacker’s intent and affects the system’s overall output. To demonstrate the IPI scenario, we set up a malicious web server containing HTML content with false instructions. This attack aimed to lure the agent into repeatedly following deceptive URLs, resulting in another form of system exhaustion. Lastly, in the MP scenario, we poisoned the RAG database that the system’s agents rely on, which results in poisoned information being retrieved.
5.1.3. Data Collection and Splitting
@@ 5.1.4. LLM Configuration
To demonstrate the robustness of our framework, we evaluate its effectiveness using two different LLMs as the agents’ underlying models: OpenAI GPT-4o mini hurst2024gpt and OpenAI o3-mini openai2025o3mini_systemcard. GPT-4o mini is a lightweight version of GPT-4o, optimized for speed and efficiency, with strong general performance. o3-mini is a compact LLM that has been shown to perform efficiently across various natural language processing tasks.
@@
Table 1. Anomaly detection results on CrewAI applications using GPT-4o-mini as the underlying model
Vulnerability Method Performance Overhead
Accuracy $\uparrow$ F1 $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ FPR $\downarrow$ Latency $\downarrow$ Token Count $\downarrow$
Hallucination Consistency of 3 0.681 ± 0.020 0.721 ± 0.015 0.826 ± 0.018 0.641 ± 0.018 0.464 ± 0.033 12.941 5312.4
Consistency of 5 0.684 ± 0.009 0.727 ± 0.008 0.842 ± 0.013 0.640 ± 0.006 0.474 ± 0.009 24.072 8848.4
LLM-as-a-judge 0.574 ± 0.010 0.332 ± 0.025 0.212 ± 0.019 0.767 ± 0.022 0.064 ± 0.005 6.215 2414.0
Agent-as-a-judge 0.586 ± 0.033 0.406 ± 0.059 0.284 ± 0.051 0.717 ± 0.073 0.112 ± 0.030 9.617 1436329.0
LogBERT 0.506 ± 0.005 0.320 ± 0.312 0.422 ± 0.509 0.575 ± 0.068 0.410 ± 0.520 0.006 -
EPI (Ours) 0.755 ± 0.015 0.722 ± 0.013 0.638 ± 0.040 0.838 ± 0.057 0.128 ± 0.059 0.007 -
Semantic (Ours) 0.716 ± 0.009 0.689 ± 0.016 0.632 ± 0.048 0.763 ± 0.034 0.200 ± 0.054 0.013 -
Combined (Ours) 0.765 ± 0.015 0.733 ± 0.016 0.646 ± 0.030 0.850 ± 0.039 0.116 ± 0.042 0.016 -
Bias toxic-bert 0.550 ± 0.000 0.683 ± 0.000 0.970 ± 0.000 0.527 ± 0.000 0.870 ± 0.000 0.251
LLM-as-a-judge 0.907 ± 0.008 0.909 ± 0.008 0.926 ± 0.011 0.892 ± 0.010 0.112 ± 0.011 5.297 2229.0
Agent-as-a-judge 0.894 ± 0.015 0.892 ± 0.015 0.876 ± 0.017 0.909 ± 0.023 0.088 ± 0.024 10.129 1253684.0
LogBERT 0.508 ± 0.027 0.640 ± 0.034 0.886 ± 0.139 0.506 ± 0.019 0.870 ± 0.169 0.008 -
EPI (Ours) 0.586 ± 0.010 0.605 ± 0.012 0.634 ± 0.017 0.578 ± 0.008 0.462 ± 0.004 0.005 -
Semantic (Ours) 0.612 ± 0.009 0.632 ± 0.013 0.666 ± 0.023 0.601 ± 0.007 0.442 ± 0.015 0.020 -
Combined (Ours) 0.664 ± 0.027 0.661 ± 0.020 0.654 ± 0.015 0.668 ± 0.033 0.326 ± 0.048 0.023 -
DPI LLM-as-a-judge 0.705 ± 0.016 0.750 ± 0.013 0.903 ± 0.015 0.644 ± 0.014 0.492 ± 0.025 7.268 4413.3
Agent-as-a-judge 0.729 ± 0.023 0.760 ± 0.021 0.885 ± 0.021 0.670 ± 0.022 0.427 ± 0.036 18.275 3682504.8
LogBERT 0.572 ± 0.072 0.665 ± 0.059 0.868 ± 0.119 0.550 ± 0.052 0.724 ± 0.251 0.005 -
EPI (Ours) 0.713 ± 0.018 0.734 ± 0.013 0.826 ± 0.011 0.666 ± 0.021 0.400 ± 0.038 0.005 -
Semantic (Ours) 0.815 ± 0.010 0.819 ± 0.011 0.865 ± 0.022 0.784 ± 0.013 0.234 ± 0.023 0.071 -
Combined (Ours) 0.802 ± 0.021 0.803 ± 0.026 0.842 ± 0.038 0.775 ± 0.021 0.238 ± 0.028 0.076 -
IPI LLM-as-a-judge 0.547 ± 0.004 0.681 ± 0.002 0.966 ± 0.009 0.526 ± 0.003 0.872 ± 0.015 17.987 6642.0
Agent-as-a-judge 0.614 ± 0.020 0.698 ± 0.013 0.890 ± 0.012 0.574 ± 0.015 0.662 ± 0.036 34.618 9895764.8
LogBERT 0.959 ± 0.022 0.960 ± 0.021 0.970 ± 0.017 0.951 ± 0.044 0.052 ± 0.049 0.007 -
EPI (Ours) 0.943 ± 0.011 0.944 ± 0.012 0.972 ± 0.029 0.919 ± 0.012 0.086 ± 0.015 0.010 -
Semantic (Ours) 0.910 ± 0.013 0.910 ± 0.012 0.912 ± 0.008 0.909 ± 0.023 0.092 ± 0.026 0.079 -
Combined (Ours) 0.941 ± 0.033 0.941 ± 0.033 0.942 ± 0.039 0.941 ± 0.040 0.060 ± 0.041 0.102 -
MP LLM-as-a-judge 0.797 ± 0.012 0.814 ± 0.011 0.890 ± 0.014 0.750 ± 0.011 0.296 ± 0.015 7.835 3701.0
Agent-as-a-judge 0.746 ± 0.013 0.749 ± 0.016 0.758 ± 0.028 0.740 ± 0.011 0.266 ± 0.015 10.484 1770939.8
LogBERT 0.734 ± 0.009 0.765 ± 0.001 0.864 ± 0.036 0.687 ± 0.023 0.396 ± 0.054 0.004 -
EPI (Ours) 0.523 ± 0.011 0.626 ± 0.019 0.800 ± 0.047 0.515 ± 0.006 0.754 ± 0.030 0.003 -
Semantic (Ours) 0.715 ± 0.019 0.714 ± 0.019 0.712 ± 0.048 0.719 ± 0.036 0.282 ± 0.065 0.027 -
Combined (Ours) 0.477 ± 0.068 0.536 ± 0.065 0.608 ± 0.097 0.483 ± 0.058 0.654 ± 0.114 0.032 -
5.1.5. Baselines
@@
As no prior work had been done to detect failures in MASs, we utilized the LLM-as-a-judge paradigm zheng2023judging with an engineered prompt for the anomaly detection task. Inspired by zhuge2024agent zhuge2024agent, we also implemented the agent-as-a-judge variant, with the task and agent definition based on the same engineered prompt. We equip the agent with scraping and searching tools, enabling it to retrieve additional information to improve its effectiveness in detecting failures. GPT-4o hurst2024gpt was used as the backbone of our baselines, analyzing each agent’s input and output. Consequently, the prediction of the baseline approaches also indicates which agent is responsible for the failure, and this information is used in our RCA evaluation. Additional results using the smaller and more cost-efficient GPT-4o-mini hurst2024gpt are provided in Appendix C. As a non-LLM baseline, we implemented LogBERT guo2021logbert for log-based anomaly detection, providing a purely sequential modeling reference against our anomaly detectors, more details are provided in Appendix C. Furthermore, we evaluate existing hallucination and bias detection strategies for our FTA apps. For hallucination detection, we adopted the consistency comparison approach manakul2023selfcheckgpt, in which each agent was instructed to repeat the task 3 or 5 additional times, and the original result was considered consistent (i.e., benign) if it aligned with the majority of the outputs. Regarding bias, we employed toxic-bert Detoxify to assess the bias in each corresponding agent.
5.1.6. Metrics
@@
To evaluate failure detection performance, we used several standard metrics applied in binary classification tasks (accuracy, precision, recall, F1 score, and false positive rate (FPR)), as well as metrics measuring the solutions’ overhead resource consumption (average inference time and token count for models that require payment). For the RCA evaluation, we used the root cause accuracy (RA) metric from zhang2024mabc; zhou2023d. In the results tables, the best performance in each column is marked in bold, and the second-best is underlined. The arrows ($\uparrow$, $\downarrow$) indicate whether the optimal value is higher or lower.
5.1.7. Implementation Details
@@
We implemented our method using Python 3.11 and PyTorch 2.6, and the models were trained on an RTX 4090 GPU. Each application had its own set of hyperparameters, and all models were trained using the Adam optimizer. For the semantic features, the textual features (i.e., the agent’s LLM interactions) were embedded using the all-MiniLM-L6-v2 model based on sentence-transformers reimers2019sentence. A bidirectional LSTM architecture was used in the AE model. The anomaly detection classification threshold was selected based on the F1 score on the validation set, while the models’ other hyperparameters were tuned to minimize the reconstruction error on the benign validation set. The distance function used in all three models, loss function was the mean square error. The exact hyperparameters are provided in Appendix C.
5.2. Evaluation Results